OpenStudy (january123):

Evaluate the expression below at x = 6. A. 20 B. 235 C. 55 D. 1,315

OpenStudy (january123):

https://media.studyisland.com/cgi-bin/mimetex.cgi? \left(\frac{1}{6}\right)x^{3}\%20+\%2019

OpenStudy (anonymous):

Don't know what you wrote, but plug in 6 where ever there is an x and evaluate.

OpenStudy (acxbox22):

lol yep the link doesnt work

OpenStudy (january123):

wait i put that ther?

OpenStudy (january123):

wrong one https://media.studyisland.com/cgi-bin/mimetex.cgi? \left(\frac{1}{6}\right)x^{3}\%20+\%2019

OpenStudy (january123):

UGHH!

OpenStudy (january123):

O WAIT NOO!! U HAVE TO COPY AND PASTE THE WHOLE THING IN A NEW TAB

OpenStudy (january123):

o sorry caps^